OpenStudy (anonymous):

what is the value of 2^2*2^3/2^-2*2^-3

OpenStudy (anonymous):

2^2+3+2+3=2^10

OpenStudy (sriram):

4*8 32 ----- =------ =32*32=1024 1/4*1/8 1/32 i guess uzma,s method is better

OpenStudy (anonymous):

16